why do you rinse the buret with the sodium hydroxide solution and not with distilled water?
how would your titration results be affected if you rinsed with distilled water before filling your buret?
1
Expert's answer
2015-11-09T08:24:10-0500
The titration is an analytical method, which allows to estimate the concentration of the given solution. Typically in this method, chemists use burettes filled with a solution of the compound which reacts with the compound of interst. Due to the fact, that all burettes are made of glass, it can absorb and remain water on the surface, because of the polarity of the glass and intermolecular forces. Sometimes chemists perform micro analysis, that is why even slightly difference in the concentration will drastically affect to the result. Thus, you have to rinse the burette with a solution which must be filled in it, because distilled water change the concentration of the initial solution.
If you rinse with water, the remaining water droplets will decrease the concentration of the solution and therefore you'll need bigger volume of the titrant, to perform the analysis. Usually the concentration can be calculated with a formula:
C1*V1=C2*V2, so C2=(C1*V1)/V2
So when water is added, we take the standard concentration of the titrant (e.g. NaOH), but bigger volume of it. So in the end you'll have the incorrect result and you'll get higher concentration than it is.